What is -x+4y+3=2x-7 written in slope-intercept form?

Answer by TimothyLamb(4379)

Get the equation into slope-intercept form: y = mx + b
by algebra like so:
---
-x + 4y + 3 = 2x - 7
4y = 2x + x - 7 - 3
4y = 3x - 10
y = (3/4)x - (10/4)
Answer:
y = (3/4)x - (5/2)
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php